OpenStudy (anonymous):

What is the equation of a line, in general form, with a point (-3, 0) and an undefined slope?

OpenStudy (anonymous):

please help this is some of my homework and I really do not get this problem

OpenStudy (anonymous):

slope is undefined so slope is \[\frac{ 1 }{ 0 }\] in order to determine the equation of line with a point (x1,y1) we will use the formulae given below y-y1 = m*(x-x1) where m= slope